# Effect of flow-aligned external magnetic fields on mushroom instability

### v1

@@ Abstract
Mushroom instability (MI) is a shear instability considered responsible for generating and amplifying magnetic fields in relativistic jets. While astrophysical jets are usually considered to be magnetized, how MI acts in magnetized jets remains poorly understood. In this paper, we investigate the effect of a flow-aligned external magnetic field on MI, with both theoretical analyses and particle-in-cell (PIC) simulations. In the limit of a cold and collisionless plasma, we derive a generalized dispersion relation for linear growth rates of the magnetized MIs. Numerical solutions of the dispersion relation reveal that the external magnetic field always suppresses the growth of MI, though MIs are much more robust to the external magnetic field than electron-scale Kelvin-Helmholtz instabilities (ESKHIs). Analyses are also extended to instabilities with an arbitrary wavevector in the shear interface plane. Two-dimensional PIC simulations of single-mode MIs reach a good agreement with our analytical predictions. In simulations with finite temperatures, we observe the competition and cooperation between MIs and a diffusion-induced DC magnetic field.
preprint: APS/123-QED
I Introduction
@@
A possibility is that the dissipation processes are (partly) mediated by shear instabilities in the neighborhood of the jet boundary, where a sharp transverse velocity gradient exists between the jet and the ambient plasma. In the collisionless regime, 3D particle-in-cell (PIC) simulations showed that the jet boundary is unstable to electromagnetic instabilities (Alves et al., 2012), where small fluctuations of electromagnetic fields are dramatically magnified by dissipating the kinetic energy of the shear flow, until being saturated. Theoretical analyses and 2D simulations conducted in separate planes have revealed that the electromagnetic fields are actually magnified by electron-scale Kelvin-Helmholtz instabilities (ESKHIs; Gruzinov (2008)), mushroom instabilities (MIs; Alves et al. (2015)) and their coupling if in 3D scenarios. Both kinds of instabilities are dominated by electron (and also positron if any) dynamics and take place on a timescale of $1/\omega_{pe}$, with ESKHIs operating in planes parallel to the velocity shear (defined by the directions of the flow velocity and its gradient) and MIs in planes perpendicular to the flow velocity. While ESKHIs are characterized with vortex structures of electrons, MIs exhibit a mushroom-like structure (Alves et al., 2015) in their nonlinear stage, resembling the morphology of classical Rayleigh-Taylor instabilities. Particularly, Alves et al. (2015) showed that while the maximum linear growth rate of MI ($\sigma_{\text{max}}=\omega_{pe}\gamma_{0}V_{0}/c$ for a symmetrical shear, where $V_{0}$ is the shear velocity and $\gamma_{0}$ is the corresponding Lorentz factor) is smaller than that of ESKHI ($\sigma_{\text{max}}=\omega_{pe}/2\sqrt{2}$) for subrelativistic shear velocities ($V_{0}<c/3$), instead it is much larger than the latter in the relativistic limit, indicating the dominance of MI in relativistic jets. Recently, Kawashima et al. (2022) proposed that MI can drive magnetic reconnection, which can trigger particle acceleration and lead to jet spine formation, and both are characteristic problems in jets. Therefore, investigating the behavior of MI is significant for us to understand the mechanism of jets.
@@
In this paper, we study the effect of a flow-aligned external magnetic field on MI with both theoretical analyses and PIC simulations. In Sec. II, we give a description of the problem and derive a generalized eigenequation for linear growth rates of MIs. Numerical solutions of the dispersion relation are also presented. In Sec. III, we present and analyze the results of our PIC simulations to verify the theoretical results. Simulations are carried out either in cold plasmas with single-mode perturbations, or in thermal plasmas with no initial perturbation. In Sec. IV, we draw a conclusion and present some discussions. The mathematical details of deriving the eigenequation are presented in Appendix A. We use SI units throughout the paper.
II Theoretical analyses
@@
II.1 Physical settings of the magnetized shear flow
@@
Now we consider a shear flow with a non-uniform velocity distribution $\textbf{{v}}_{e0}=\textbf{{v}}_{i}=[0,v_{0}(x),0]$. Still, the ions are considered to be free-streaming and unperturbed throughout the timescale we investigate. The initial number density distribution is $n_{e0}(x)=n_{i}(x)$, maintaining the charge and current neutrality of the shear flow. A flow-aligned external magnetic field $\textbf{{B}}_{0}=[0,B_{0}(x),0]$ is applied to the system. In the cold plasma with thermal pressure vanishing, the dynamical equilibrium of ions and electrons requires $\textbf{{E}}_{0}=0$, given $\textbf{{v}}_{e0}\times\textbf{{B}}_{0}=0$. No net current $\textbf{{J}}_{0}$ results in a uniform magnetic field $B_{0}(x)=\text{const.}$ We note that uniform external magnetic fields with components in all three directions were considered by Che (2025) for ESKHIs in incompressible plasmas; However, this situation will bring much complexity in our sets of equations, leading us to start with the simple case of a flow-aligned field.
II.2 The eigenequation and the dispersion relation of MI linear growth rate
@@
where $G(k_{z})=(1+k_{z}^{2}c^{2}/\omega_{p}^{2})$. This equation predicts the maximum growth rate at an infinite wavenumber: $\sigma_{\text{max}}=\sigma(k_{z}\xrightarrow[]{}\infty)=\gamma_{0}V_{0}\omega_{p}/c$. On contrast, when $\omega_{c}\neq 0$, the equations are highly coupled and have to be solved numerically. When the shear flow is uniform in each half of the plane, one can assume solutions with the form of $\tilde{E}_{y}^{\pm}=A_{1}^{\pm}\exp(a^{\pm}x)+A_{2}^{\pm}\exp(b^{\pm}x)$ and $\tilde{E}_{z}^{\pm}=B_{1}^{\pm}\exp(a^{\pm}x)+B_{2}^{\pm}\exp(b^{\pm}x)$ $F(x)=F^{+}$ at $x>0$ and $F(x)=F^{-}$ at $x<0$ respectively], where the real part of $a^{+}$ and $b^{+}$ are restricted to be negative, and that of $a^{-}$ and $b^{-}$ positive, guaranteeing a mode vanishing at infinities. Thus, Eqs. 6 and 7 are transformed into polynomial equations of $A_{i}^{\pm},B_{i}^{\pm},a^{\pm}$$b^{\pm}$, and also $\omega$. Along with the boundary conditions of the electric field, the equations can be solved straightforward to give $\omega$ (see the Appendix in Guo et al. (2025) for detail).
We numerically solve MI growth rates $\sigma$ for symmetrical shear flows $n_{e0}(x)=const.,v_{0}(x)=\text{sgn}(x)V_{0},$ and $\omega_{c}(x)=eB_{0}/\gamma_{0}m_{e}$] with certain values of $V_{0}$ and $\omega_{c}/\omega_{p}$. Figs. 1, 2, and 3 show the dispersion relation $\sigma(k_{z})$ for $V_{0}/c=0.2$, 0.5 and 0.8 respectively. Especially, growth rates extracted from single-mode PIC simulations are also marked in Fig. 2, which will be discussed in detail in Sec. III.
@@
Figure 2: The same plot as Fig. 1, but with $V_{0}=0.5c$. Results extracted from single-mode PIC simulations are also marked for comparison, where the horizontal coordinate of the marks corresponds to the wavenumber of the pre-imposed perturbation, and the vertical coordinate corresponds to the growth rates extracted from the simulations. The details of the simulations will be discussed in Sec. III.
@@
II.3 Competition between MI and ESKHI
@@
Through different values of $V_{0}$, it is obvious that MIs are more robust to the external magnetic field than ESKHIs. For each value of $V_{0}$, ESKHIs are completely stabilized by external magnetic fields larger than a threshold value. On contrast, MIs remain unstable under large values of $\omega_{c0}/\omega_{p0}$. In consequence, shear flows with $V_{0}/c<1/3$ should be dominated by ESKHIs under weak external magnetic fields, but dominated by MIs under strong external magnetic fields (see the crossing of the solid and dashed blues curves for $V_{0}=0.2c$ in Fig. 4). Our results also imply that a finite threshold stabilizing magnetic field may not exist for MIs in the cold plasma limit.
@@
II.4 The dispersion relation of the coupled instability
@@
Figure 5 shows the numerical solution of $\sigma(k_{y},k_{z})$ in a symmetrical shear flow with $V_{0}=0.2c$. The dispersion relation of the coupled instability inherits some characteristics from MI and ESKHI. Like MI, the unstable region of the coupled instability extends to an infinite $k_{z}$; On the other hand, the unstable region is cut off at some $k_{y}$, resembling the dispersion relation of ESKHI. In the unmagnetized scenario with $\omega_{c}=0$ (the blue surface), the dispersion relation basically looks like a sweeping of ESKHI dispersion relation (the yellow curve at $k_{z}=0$) over $k_{z}$, except that the maximum growth rate slightly increased with $k_{z}$. Interestingly, under the external magnetic field $\omega_{c}=0.5\omega_{p}$ (the red surface), the dispersion relation is distorted towards the $k_{y}=0$ axis as $k_{z}$ increases. Especially, the region with $0.05\lesssim k_{y}V_{0}/\omega_{p}\lesssim 0.4$ and $k_{z}V_{0}/\omega_{p}\gtrsim 0.4$ is destabilized in this case (where the red surface is above the blue one), while both ESKHIs and MIs in separate are stabilized by this magnetic field.
@@ III PIC simulations
III.1 Single-mode simulations
@@
We also care about the growth rate of the instability in simulations for giving a quantitative verification for our analytical results in Sec. II. Figure 8 shows the evolution of the generated magnetic energy $\Delta E_{B}=\int dV(B-B_{0})^{2}/2\mu_{0}$ in different runs. In each run, a linear growth stage of the instability followed by a saturation stage is clearly observed. Theoretically, in the linear stage of the instability there is $\Delta E_{B}\propto\delta B^{2}\propto\exp(2\sigma t)$, so we evaluate the linear growth rate $\Gamma_{s}$ of $\Delta E_{B}$ in each run, and take $\sigma_{s}=\Gamma_{s}/2$ as the growth rate of MIs observed in simulations. The results of $\sigma_{s}$ are marked at corresponding wavenumbers $k_{z}$ in Fig. 2, where all the simulation results of $\sigma_{s}$ are close to the analytical dispersion relation of $\sigma(k_{z})$. This is not surprising, as in a cold and collisionless plasma our equations in Sec. II.1 give a sufficiently accurate description of the plasma dynamics.
III.2 Simulations with finite temperatures
@@
In simulations with higher temperatures, a more different phenomenon can be resolved. Figure 10 shows $J_{ey}$ in the $T=100$ eV run with other settings the same. In Panel (a), on each side of the shear interface, one can observe a DC component of $J_{ey}$ in a direction opposite to the bulk current. This results from thermal diffusion of electrons across the shear interface, which cannot be described by our fluid theory but requires a kinetic approach. While ions remain stationary, the electrons crossing the interface bring a $v_{y}$ in the opposite direction, leading to a current imbalance on each side of the interface. Consequently, a DC component of the magnetic field $B_{\text{DC}}\sim 4\sqrt{2\pi}en_{0}\beta_{0}v_{the}t$ is generated in the $z$ direction (Grismayer et al., 2013), which shares the same direction with the DC magnetic field formed in the saturation stage of MI [see Fig. 6 (d) and (h)]. Since the AC magnetic field of MI depends on time as $\delta\textbf{{B}}\sim\textbf{{B}}_{f}\exp(\sigma t)$ and has to grow from small thermal fluctuations $\textit{{B}}_{f}$, $B_{\text{DC}}$ will dominate in the beginning stage with the $v_{the}t$ dependence. Afterwards, the exponential dependence of MI modes will take over, forming mushroom structure from fluctuations [Panel (b)], but tainted with DC filaments [Panel (c)]. When MI enters the saturation stage, its nonlinear field appears again in the DC manner, merging into the DC magnetic field induced by thermal diffusion.
@@
The competition and cooperation between the diffusion-induced DC magnetic field and the MI modes is also pronounced in the evolution of the magnetic energy. Figure 11 shows the evolution of $\Delta E_{B}(t)/E_{p}(0)$ in the thermal runs. As seen in Panel (a), in the 1 and 10 eV runs the growth stage starts with a sharp increase that gradually smooths out, corresponding to the $v_{the}t$ dependence of the diffusion-induced DC magnetic field (note the logarithmic y-axis of the figure). Then, a linear growth stage corresponding to the MI magnetic field $B_{f}\exp(\sigma t)$ follows, where the order of $\Delta E_{B}$ increases the most. On contrary, in the 100 and 1000 eV runs, the diffusion-induced field grows faster and larger, and affects the whole growth stage more, and the linear growth stage of MI can hardly be resolved from the $\Delta E_{B}$ evolution. However, at a late time, $\Delta E_{B}(t)/E_{p}(0)$ in all runs reaches about $3\times 10^{-4}$ whatever the temperature, which indicates that the diffusion-induced DC field and the MI field share a similar nature. To see this directly, one can imagine superposing a perturbation in an opposite direction in Fig. 6(a). This effectively gives the plasma a thermal velocity in the $x$ direction and allows the two bulk to diffuse into each other. We note that in Fig. 6(a), the magnetic fields are displaced from the initial shear interface in the $x$ direction. In consequence, the superposition of an opposite mode will not cancel the fields (while in linearized fluid theory, such superposition will only lead to cancellation). Instead, their combination will appear as a DC mode of $B_{z}$ along the shear interface, similar to that caused by thermal diffusion of electrons. This qualitatively explains the relation between the MI fields and the thermal-induced DC fields.
@@
As seen in Fig. 11, in the linear stage of each of the thermal runs, we observe a growth rate smaller than the MI maximum growth rate predicted by our cold fluid theory. In particular, as temperature increases, the growth rate (if there is a linear stage) even slightly decreases in the unmagnetized runs [Fig. 11(a)]. This is contrary to the conclusion in Miller & Rogers (2016) that the MI growth rate increases as temperature becomes larger. The reason is that the former work still adopted the fluid approach which neglects the interfacial diffusion. Although higher temperatures truly result in a larger thermal pressure and a faster dynamical response of the plasmas, in our simulations the diffusion-induced magnetic field partly magnetizes the shear interface in advance and pushes the electrons away, therefore suppressing the subsequent linear growth of MI (Guo et al., 2025).

### v2

@@ Abstract
Mushroom instability (MI) is a shear instability considered responsible for generating and amplifying magnetic fields in relativistic jets. While astrophysical jets are usually magnetized, how MI acts in magnetized jets remains poorly understood. In this paper, we investigate the effect of a flow-aligned external magnetic field on MI, with both theoretical analyses and particle-in-cell (PIC) simulations. In the limit of a cold and collisionless plasma, we derive a generalized dispersion relation for linear growth rates of the magnetized MIs. Numerical solutions of the dispersion relation reveal that the external magnetic field always suppresses the growth of MI, though MIs are much more robust against the external magnetic field than electron-scale Kelvin-Helmholtz instabilities (ESKHIs). Analyses are also extended to instabilities with an arbitrary wavevector in the shear interface plane, where coupling effect is observed for sub-relativistic scenarios. Two-dimensional PIC simulations of single-mode MIs reach a good agreement with our analytical predictions, and we observe formation of a quasi-steady saturation structure in magnetized runs. In simulations with finite temperatures, we observe the competition and cooperation between MIs and a diffusion-induced DC magnetic field.
Relativistic jets(1390) — Plasma jets(1263) — Plasma physics(2089) — Magnetic fields(994)
1 Introduction
@@
A possibility is that the dissipation processes are (partly) mediated by shear instabilities in the neighborhood of the jet boundary, where a sharp transverse velocity gradient exists between the jet and the ambient plasma. In the collisionless regime, 3D particle-in-cell (PIC) simulations showed that the jet boundary is unstable to electromagnetic instabilities (Alves et al., 2012), where small fluctuations of electromagnetic fields are dramatically magnified by dissipating the kinetic energy of the shear flow, until being saturated. Theoretical analyses and 2D simulations conducted in separate planes have revealed that the electromagnetic fields are actually magnified by electron-scale Kelvin-Helmholtz instabilities (ESKHIs; Gruzinov (2008)), mushroom instabilities (MIs; Alves et al. (2015)) and their coupling if in 3D scenarios. Both kinds of instabilities are dominated by electron (and also positron, if present) dynamics and take place on a timescale of $1/\omega_{pe}$. While ESKHIs operate in the plane of the shear flow (with a wavevector parallel to the flow direction), MIs develop in planes perpendicular to the flow velocity (with a wavevector perpendicular to the shear plane). While ESKHIs are characterized with vortex structures of electrons, MIs exhibit a mushroom-like structure (Alves et al., 2015) in their nonlinear stage, resembling the morphology of classical Rayleigh-Taylor instabilities. Particularly, Alves et al. (2015) showed that while the maximum linear growth rate of MI ($\sigma_{\text{max}}=\omega_{pe}\gamma_{0}V_{0}/c$ for a symmetrical shear, where $V_{0}$ is the shear velocity and $\gamma_{0}$ is the corresponding Lorentz factor) is smaller than that of ESKHI ($\sigma_{\text{max}}=\omega_{pe}/2\sqrt{2}$) for subrelativistic shear velocities ($V_{0}<c/3$), instead it is much larger than the latter in the relativistic limit, indicating the dominance of MI in relativistic jets. Recently, Kawashima et al. (2022) proposed that MI can drive magnetic reconnection, which can trigger particle acceleration and lead to jet spine formation, and both are characteristic problems in jets. Therefore, investigating the behavior of MI is significant for us to understand the mechanism of jets.
@@
In this paper, we study the effect of a flow-aligned external magnetic field on MI with both theoretical analyses and PIC simulations. In Sec. 2, we give a description of the problem and derive a generalized eigenequation for linear growth rates of MIs. Numerical solutions of the dispersion relation are also presented. In Sec. 3, we present and analyze the results of our PIC simulations to verify the theoretical results. Simulations are carried out either in cold plasmas with single-mode perturbations, or in thermal plasmas with no initial perturbation. In Sec. 4, we draw a conclusion and present some discussions. The mathematical details of deriving the eigenequation are presented in Appendix A. We use SI units throughout the paper.
2 Theoretical analyses
@@
2.1 Physical settings of the magnetized shear flow
@@
Now we consider a shear flow with a non-uniform velocity distribution $\textbf{{v}}_{e0}=\textbf{{v}}_{i}=[0,v_{0}(x),0]$. Still, the ions are considered to be free-streaming and unperturbed throughout the timescale we investigate. The initial number density distribution is $n_{e0}(x)=n_{i}(x)$, maintaining the charge and current neutrality of the shear flow. A flow-aligned external magnetic field $\textbf{{B}}_{0}=[0,B_{0}(x),0]$ is applied to the system. In the cold plasma with thermal pressure vanishing, the dynamical equilibrium of ions and electrons requires $\textbf{{E}}_{0}=0$, given $\textbf{{v}}_{e0}\times\textbf{{B}}_{0}=0$. No net current $\textbf{{J}}_{0}$ results in a uniform magnetic field $B_{0}(x)=\text{const.}$ We note that uniform external magnetic fields with components in all three directions were considered by Che (2025) for ESKHIs in incompressible plasmas; However, this situation will bring much complexity and especially discontinuities in our sets of equations, leading us to start with the simple case of a flow-aligned field.
2.2 The eigenequation and the dispersion relation of MI linear growth rate
@@
where $G(k_{z})=(1+k_{z}^{2}c^{2}/\omega_{p}^{2})$. This equation predicts the maximum growth rate at an infinite wavenumber: $\sigma_{\text{max}}=\sigma(k_{z}\xrightarrow[]{}\infty)=\gamma_{0}V_{0}\omega_{p}/c$. In contrast, when $\omega_{c}\neq 0$, the equations are highly coupled and have to be solved numerically. When the shear flow is uniform in each half of the plane, one can assume solutions with the form of $\tilde{E}_{y}^{\pm}=A_{1}^{\pm}\exp(a^{\pm}x)+A_{2}^{\pm}\exp(b^{\pm}x)$ and $\tilde{E}_{z}^{\pm}=B_{1}^{\pm}\exp(a^{\pm}x)+B_{2}^{\pm}\exp(b^{\pm}x)$ $F(x)=F^{+}$ at $x>0$ and $F(x)=F^{-}$ at $x<0$ respectively], where the real part of $a^{+}$ and $b^{+}$ are restricted to be negative, and that of $a^{-}$ and $b^{-}$ positive, guaranteeing a mode vanishing at infinities. Thus, Eqs. 6 and 7 are transformed into polynomial equations of $A_{i}^{\pm},B_{i}^{\pm},a^{\pm}$$b^{\pm}$, and also $\omega$. Along with the boundary conditions of the electric field, the equations can be solved straightforward to give $\omega$ (see the Appendix in Guo et al. (2025) for detail).
We numerically solve MI growth rates $\sigma$ for symmetrical shear flows $n_{e0}(x)=const.,v_{0}(x)=\text{sgn}(x)V_{0},$ and $\omega_{c}(x)=eB_{0}/\gamma_{0}m_{e}$] with certain values of $V_{0}$ and $\omega_{c}/\omega_{p}$. Figs. 1, 2, and 3 show the dispersion relation $\sigma(k_{z})$ for $V_{0}/c=0.2$, 0.5 and 0.8 respectively. Especially, growth rates extracted from single-mode PIC simulations are also marked in Fig. 2, which will be discussed in detail in Sec. 3.
@@
Figure 2: The same plot as Fig. 1, but with $V_{0}=0.5c$. Results extracted from single-mode PIC simulations are also marked for comparison, where the horizontal coordinate of the marks corresponds to the wavenumber of the pre-imposed perturbation, and the vertical coordinate corresponds to the growth rates extracted from the simulations. The details of the simulations will be discussed in Sec. 3.
@@
2.3 Competition between MI and ESKHI
@@
For all values of $V_{0}$ considered, it is obvious that MIs are more robust against the external magnetic field than ESKHIs. For each value of $V_{0}$, ESKHIs are completely stabilized by external magnetic fields larger than a threshold value. In contrast, MIs remain unstable under large values of $\omega_{c0}/\omega_{p0}$. In consequence, shear flows with $V_{0}/c<1/3$ should be dominated by ESKHIs under weak external magnetic fields, but dominated by MIs under strong external magnetic fields (see the crossing of the solid and dashed blues curves for $V_{0}=0.2c$ in Fig. 4). Our results also imply that a finite threshold stabilizing magnetic field may not exist for MIs in the cold plasma limit.
@@
2.4 The dispersion relation of the coupled instability
@@
Figure 5 shows the numerical solution of $\sigma(k_{y},k_{z})$ in a symmetrical shear flow with $V_{0}=0.2c$. The dispersion relation of the coupled instability inherits some characteristics from MI and ESKHI. Like MI, the unstable region of the coupled instability extends to an infinite $k_{z}$; On the other hand, the unstable region is cut off at some $k_{y}$, which is the same as the dispersion relation of ESKHI. In the unmagnetized case with $\omega_{c}=0$ (the blue surface), the dispersion relation largely resembles that of ESKHI dispersion relation (the yellow curve at $k_{z}=0$) extended over $k_{z}$, with the maximum growth rate slightly increasing with $k_{z}$. Interestingly, under the external magnetic field $\omega_{c}=0.5\omega_{p}$ (the red surface), the dispersion relation is distorted towards the $k_{y}=0$ axis as $k_{z}$ increases. Especially, the region with $0.05\lesssim k_{y}V_{0}/\omega_{p}\lesssim 0.4$ and $k_{z}V_{0}/\omega_{p}\gtrsim 0.4$ is destabilized in this case (where the red surface is above the blue one), while both ESKHIs and MIs in separate are stabilized by this magnetic field.
@@ 3 PIC simulations
3.1 Single-mode simulations
@@
We also care about the growth rate of the instability in simulations for giving a quantitative verification for our analytical results in Sec. 2. Figure 8 shows the evolution of the generated magnetic energy $\Delta E_{B}=\int dV(B-B_{0})^{2}/2\mu_{0}$ in different runs. In each run, a linear growth stage of the instability followed by a saturation stage is clearly observed. Theoretically, in the linear stage of the instability there is $\Delta E_{B}\propto\delta B^{2}\propto\exp(2\sigma t)$, so we evaluate the linear growth rate $\Gamma_{s}$ of $\Delta E_{B}$ in each run, and take $\sigma_{s}=\Gamma_{s}/2$ as the growth rate of MIs observed in simulations. The results of $\sigma_{s}$ are marked at corresponding wavenumbers $k_{z}$ in Fig. 2, where all the simulation results of $\sigma_{s}$ are close to the analytical dispersion relation of $\sigma(k_{z})$. This is not surprising, as in a cold and collisionless plasma our equations in Sec. 2.1 give a sufficiently accurate description of the plasma dynamics.
3.2 Simulations with finite temperatures
@@
In simulations with higher temperatures, a more different phenomenon can be resolved. Figure 10 shows $J_{ey}$ in the $T=100$ eV run with other settings the same. In Panel (a), on each side of the shear interface, one can observe a DC component of $J_{ey}$ in a direction opposite to the bulk current. This results from thermal diffusion of electrons across the shear interface, which cannot be described by our fluid theory but requires a kinetic approach. While ions remain stationary, the electrons crossing the interface bring a $v_{y}$ in the opposite direction, leading to a current imbalance on each side of the interface. Consequently, a DC component of the magnetic field $B_{\text{DC}}\sim 4\sqrt{2\pi}en_{0}\beta_{0}v_{the}t$ is generated in the $z$ direction (Grismayer et al., 2013), which shares the same direction with the DC magnetic field formed in the saturation stage of MI [see Fig. 6 (d) and (h)]. Since the AC magnetic field of MI, which depends on time as $\delta\textbf{{B}}\sim\textbf{{B}}_{f}\exp(\sigma t)$, must grow from small thermal fluctuations $\textit{{B}}_{f}$, $B_{\text{DC}}$ will dominate at early times with the $v_{the}t$ dependence. Afterwards, the exponential dependence of MI modes will take over, forming mushroom structure from fluctuations [Panel (b)], but tainted with DC filaments [Panel (c)]. When MI enters the saturation stage, its nonlinear field appears again in the DC manner, merging into the DC magnetic field induced by thermal diffusion.
@@
The competition and cooperation between the diffusion-induced DC magnetic field and the MI modes is also pronounced in the evolution of the magnetic energy. Figure 11 shows the evolution of $\Delta E_{B}(t)/E_{p}(0)$ in the thermal runs. As seen in Panel (a), in the 1 and 10 eV runs the growth stage starts with a sharp increase that gradually smooths out, corresponding to the $v_{the}t$ dependence of the diffusion-induced DC magnetic field (note the logarithmic y-axis of the figure). Then, a linear growth stage corresponding to the MI magnetic field $B_{f}\exp(\sigma t)$ follows, where the order of $\Delta E_{B}$ increases the most. In contrast, in the 100 and 1000 eV runs, the diffusion-induced field grows faster and larger, and affects the whole growth stage more, and the linear growth stage of MI can hardly be resolved from the $\Delta E_{B}$ evolution. However, at a late time, $\Delta E_{B}(t)/E_{p}(0)$ in all runs reaches about $3\times 10^{-4}$ whatever the temperature, which indicates that the diffusion-induced DC field and the MI field share a similar nature. To see this directly, one can imagine superposing a perturbation in an opposite direction in Fig. 6(a). This effectively gives the plasma a thermal velocity in the $x$ direction and allows the two bulk to diffuse into each other. We note that in Fig. 6(a), the magnetic fields are displaced from the initial shear interface in the $x$ direction. In consequence, the superposition of an opposite mode will not cancel the fields (while in linearized fluid theory, such superposition will only lead to cancellation). Instead, their combination will appear as a DC mode of $B_{z}$ along the shear interface, similar to that caused by thermal diffusion of electrons. This qualitatively explains the relation between the MI fields and the thermal-induced DC fields.
@@
As seen in Fig. 11, in the linear stage of each of the thermal runs, we observe a growth rate smaller than the MI maximum growth rate predicted by our cold fluid theory. In particular, as temperature increases, the growth rate (if there is a linear stage) even slightly decreases in the unmagnetized runs [Fig. 11(a)]. This is contrary to the conclusion in Miller & Rogers (2016) that the MI growth rate increases as temperature becomes larger. This discrepancy arises because the earlier study employed a fluid approach that neglects interfacial diffusion of electrons. Although higher temperatures truly result in a larger thermal pressure and a faster dynamical response of the plasmas, in our simulations the diffusion-induced magnetic field partly magnetizes the shear interface in advance and pushes the electrons away, therefore suppressing the subsequent linear growth of MI (Guo et al., 2025).